Casey walks from A to B at 3mph and returns by bicycle at 12mph. Her return trip takes 2 hours less than her going trip. Find out how long it took to walk. Find out how long the trip took by bicycle. Find out how long the entire trip took. Find out how far it is from A to B
Answer by checkley71(8403)

DISTANCE=3 X HOURS+2 HOURS & D=12 X HOURS
DISTANCE IS THE SAME OR 3H+2=12H OR 2=9H OR H=2/9
PROOF 3(2/9)+2=12(2/9) OR (6/9)+2=24/9 OR (6/9)+(18/9=24/9 OR (6+18)/9=24/9
OR 24/9=24/9.